OpenStudy (anonymous):

Write each expression in terms of sine and cosine and simplify so that no quotients appear in the final expression. cot x sin x

OpenStudy (anonymous):

replace \(\cot(x)\) by \(\frac{\cos(x)}{\sin(x)}\)

OpenStudy (anonymous):

after I replace cot, what would the next step be ?

OpenStudy (anonymous):

(cos x*sin x)/sin x . Simplify.

jimthompson5910 (jim_thompson5910):

As satellite73 is saying you start off with this expression